someone help me i need to figure out these answer can you guys show steps.

Use the substitution method to solve the system
y=x^2
y= 7x- 6
i need to find x1 x2 y1 and y2

Answer by ewatrrr(24785)

 
Hi,
Use the substitution method to solve the system
y=x^2
y= 7x- 6
x^2 = 7x-6
x^2 - 7x + 6 = 0
(x -6)(x - 1) = 0
x = 6 or x = 1, Pts of Intersection will be (1,1) and (6,36)